Grade upto college level Physical Chemistry

Calculate the volume occupied by 5.0 g of acetylene gas at 50oC and 740 mm pressure.

Amit Saxena
12 Years agoGrade upto college level

1 Answer

Navjyot Kalra
12 Years ago
Applying the general gas equation
PV = nRT = m/M RT
Here, Mol. wt. of acetylene i.e., C2H2 (M) = 26, P = 740/760 atm,
T = 50oC = 50 + 273 = 323 K
V = mRT/MP or V 5 * 0.082 *323 *760/26 *740 = 5.23 l